### v1

# C2SaferRust: Transforming C Projects into Safer Rust with NeuroSymbolic Techniques

Abstract
In recent years, there has been a lot of interest in converting C code to Rust, to benefit from the memory and thread safety guarantees of Rust. C2Rust is a rule-based system that can automatically convert C code to functionally identical Rust, but the Rust code that it produces is non-idiomatic, i.e., makes extensive use of unsafe Rust, a subset of the language that doesn’t have memory or thread safety guarantees. At the other end of the spectrum are LLMs, which produce idiomatic Rust code, but these have the potential to make mistakes and are constrained in the length of code they can process. In this paper, we present C2SaferRust, a novel approach to translate C to Rust that combines the strengths of C2Rust and LLMs. We first use C2Rust to convert C code to non-idiomatic, unsafe Rust. We then decompose the unsafe Rust code into slices that can be individually translated to safer Rust by an LLM. After processing each slice, we run end-to-end test cases to verify that the code still functions as expected. We also contribute a benchmark of 7 real-world programs, translated from C to unsafe Rust using C2Rust. Each of these programs also comes with end-to-end test cases. On this benchmark, we are able to reduce the number of raw pointers by up to 38%, and reduce the amount of unsafe code by up to 28%, indicating an increase in safety. The resulting programs still pass all test cases. C2SaferRust also shows convincing gains in performance against two previous techniques for making Rust code safer.
1 Introduction
The Rust language has emerged as a popular alternative to C, thanks to its strong memory safety and thread safety guarantees. Rust uses a memory model based on ownership and borrowing, which prevents accesses of a memory location after it has been de-allocated or freed. It also disallows mutable aliasing, which prevents data races in multi-threaded programs. At the same time, Rust has comparable runtime performance to C, making it suitable for performance-critical applications like device drivers and embedded systems.
In contrast, C code can contain both memory safety and thread safety errors, which makes it prone to security vulnerabilities. For all of these reasons, there has been a lot of interest in converting all C code to Rust. However, doing this manually is challenging and requires tremendous manpower and expertise. Just a single project, the Linux kernel, contains more than 7 million lines of C code, spread over more than 16,000 files. Thus, there is a need to develop tools to automatically convert C to Rust.
Limitations of Existing Approaches: Existing approaches for code translation fall into two broad categories. The first is transpilers or cross-compilers, which are rule-based systems. For instance, the C2Rust project [10] is a transpiler that converts C code to Rust. Although the Rust code produced by C2Rust is almost always correct, it is unsafe Rust, a subset of the language that lacks the memory and thread safety guarantees usually associated with Rust. Further, this code is non-idiomatic, i.e.,, very different from the kind of Rust code usually written by humans, making it hard to read and maintain.
There have been a few approaches [8, 30, 11, 12] that augment the transpiler framework with specific rule-based transformations, such as converting certain datatypes to other types. However, these are limited in scope, and cannot “zoom out” to transform the code at a higher granularity.
The second category of approaches is using Large Language Models (LLMs). These models are trained on vast amounts of human-written code, and thus produce code that is idiomatic, i.e., uses the most conventional idioms and constructs of a programming language. However, this code frequently has errors and is unreliable. This problem is compounded because LLM-generated code has the appearance of correctness, which makes any errors harder to spot.
Another challenge when applying LLMs for code translation is the length of input and/or output code. LLMs operate on a fixed context window, which means that they can only process inputs and generate outputs that are less than a certain predefined length. A real-world code project with dozens of files and hundreds of lines per file may not fit within this window. Further, even if the entire project does fit, it is known that LLMs tend to produce more errors as the length of the input or output increases. For these reasons, most LLM-based code translation research has focused on programs that are relatively short standalone files, such as those used in competitive programming contests.
Our Approach: In this paper, we combine the strengths of the transpiler approach and the LLM approach to design C2SaferRust- an algorithm to convert full real-world C projects into Rust. Tackling an entire codebase at once can lead to inefficiencies and increased complexity. Instead, we propose breaking down the translation process into smaller, manageable components, such as individual methods or functions. These smaller units can be translated first, and once translated, the changes can be progressively integrated into the larger project by carefully propagating the altered parameters, variables, and interfaces to the rest of the codebase. This requires strategic planning to ensure that dependencies between components are maintained and that changes flow seamlessly throughout the system.
C2SaferRust uses the C2Rust transpiler in the first step to convert all the C code to non-idiomatic unsafe Rust. Then, we incrementally convert slices of this code to safe Rust using an LLM. The transpiled unsafe Rust code thus acts as a bridge between the original C code and the LLM-generated safe Rust code. We traverse the parse tree of the transpiled code to decompose it into smaller units which can be individually translated.
We use static program analysis to analyze dependencies between these units and arrange them in an optimal order for translation. Then we augment each unit with context information, again derived using static analysis, and prompt an LLM with this information to translate the code into safe Rust. We ensure that each unit is translated correctly by compiling the code and running end-to-end test cases each time an LLM produces a translation. If there are errors, we provide the compiler or test feedback to the LLM and iteratively refine the translations.
Figure 1: An overview of our system, C2SaferRust, to translate a C program to safer Rust. We first run the C2Rust transpiler to get non-idiomatic Rust, then decompose the program into chunks. For each chunk, we slice the program to extract relevant context information and dataflow information, and compose this into a translation prompt for an LLM. The LLM produces safer Rust, which we swap back in and test end-to-end. If the program doesn’t compile or pass test cases, we give the LLM feedback and repeat.
Summary of Results: We evaluate our tool on two datasets - one contributed by us, and one from a prior work, Laertes [8].
  • On our dataset, C2SaferRust is able to reduce the amount of unsafe code by up to 28%. We are also able to substantially reduce the occurrence of raw pointers, a datatype that is exclusively associated with unsafe Rust.
  • On the Laertes dataset, C2SaferRust outperforms two selected baseline approaches on a majority of the programs for all of our metrics.
  • We show through an ablation study that each of the components of our design contributes to its overall performance.
Our Contributions: In summary, our contributions in this paper as as follows:
  1. 1.
    We adopt a novel approach of using C2Rust transpiled code as a bridge between C code and LLM-generated safe Rust, combining the strengths of rule-based transpiler approaches and LLMs. We traverse the program parse tree to decompose a large program into smaller units that can be individually translated by an LLM. We then augment each unit with context information derived from static program analysis. We test and validate the LLM-generated translations, and use feedback-guided repair to iteratively refine them.
  2. 2.
    We contribute a benchmark dataset of 7 programs consisting of utilities from the GNU Coreutils project. Each program comes with end-to-end system-level test cases, enabling us to evaluate correctness.
  3. 3.
    The ecosystem of static analysis tools for Rust is rather limited, and so we had to implement most of our analyses as Rust compiler passes, directly using the Rust compiler’s intermediate representations. This lays the groundwork to build future Rust static analysis tools. For instance, we build program slicing and callgraph analysis algorithms, which are quintessential static analysis techniques [14]. We make all of this code publicly available on GitHub [^1] https://github.com/vikramnitin9/c2saferrust.
2 Background
2.1 Safety and Unsafety in Rust
Rust is a systems language providing a safe alternative to C and C++ (which are susceptible to bugs and security vulnerabilities [7, 15, 23]) with a robust ownership system ensuring memory and thread safety without requiring garbage collection.
In order to adhere to these safety guarantees, every value in the memory is tied to the variable that owns it, and the memory is automatically de-allocated when the owning variable is no longer in scope. While ownership is exclusive to a variable, Rust also permits other variable to temporarily borrow the values. However, each borrow is further associated with a lifetime while the Rust compiler tracks their corresponding constraints. For instance, it is idiomatic for functions to borrow their arguments from their caller (for the lifetime of the function call) before restoring ownership back to the caller after the call terminates.
Rust further introduces and enforces mutable borrows and immutable borrows which are governed by the principle of exclusivity to ensure memory safety and prevent data races. An immutable borrow (&T) allows read-only access to a value and permits multiple immutable borrows to coexist simultaneously, as they do not interfere with each other. However, while immutable borrows are active, no mutable borrow of the same value is allowed. In contrast, a mutable borrow (&mut T) provides exclusive, writable access to a value, meaning no other borrows—neither mutable nor immutable—can exist simultaneously. This exclusivity ensures that data cannot be read or modified concurrently, maintaining consistency and eliminating potential race conditions. At compile time, a borrow checker enforces these rules, guaranteeing safe and predictable behavior in all borrowing scenarios [27].
The strong safety guarantees in Rust come with several challenges, namely: a) difficulty in implementing complex data structures, such as graphs, where cyclic references or multiple ownerships are common, due to strict ownership and borrowing rules; b) performance bottlenecks such as overhead from bounds checks on array access or runtime borrow checks for dynamically mutable structures; and c) complexities in integrating with system libraries or for writing low-level kernel or driver code via foreign function interface (FFI), which lack Rust’s safety guarantees [3]. To overcome these limitations, which may be too restrictive for several scenarios, Rust provides the unsafe keyword to enable developers to perform operations that the compiler may otherwise not be able to guarantee as being safe. Unsafe blocks (and functions) allow actions such as dereferencing raw pointers, accessing mutable global variables, calling unsafe functions (e.g., through FFI or compiler intrinsics), among other options [26]. While these features are crucial for tasks like interfacing with system libraries, optimizing performance-critical code, or managing complex data structures, they require developers to take full responsibility for ensuring memory safety. Developers who use unsafe must carefully design and test their code to prove its safety, as the compiler’s guarantees no longer apply in these contexts.
The ability to strategically employ unsafe blocks and/or unsafe functions accompanied by strong safety guarantees makes Rust a compelling choice for systems programming. Consequently, there has been a concerted effort to migrate legacy codebases written in C and C++ Rust, leveraging its safety features while eliminating sources of common vulnerabilities that arise due to manual memory management and unchecked behavior [23]. However, this migration process is not straightforward. C and C++ both rely on a very permissive memory models with use of raw pointers, unchecked array indexing, and general lack of enforced ownership or borrowing rules, all of which make manual translation a very arduous and error prone process [16].
2.2 Translating C to Rust using Transpilation
One of the most direct approaches to translate C code to Rust is to make use of source-to-source compilers (also known as Transpilers). Transpilers aim to translate code written in one programming language to another while endeavoring to preserve the semantic equivalence between the two. The resulting code is intended to be functionally identical to the original. In the context of translating from C to Rust, transpilers such as C2Rust [10] use carefully designed rules to map C constructs into their Rust counterparts.
Challenges with transpilers.  Source to source compilers, such as C2Rust [10], provide an automated approach to translate C to functionally equivalent Rust. However, as discussed previously, transpilers rely on a predetermined set of rule to map constructs like raw pointers, global variables, and manual memory management into their corresponding Rust equivalents. However, these mappings often result in code that compiles in Rust but compromises its safety guarantees, requiring the liberal use of unsafe blocks or foreign functions to maintain functional equivalence. This highlights the difficulty in transforming C’s permissive programming model to that of Rust’s strict ownership and borrowing models. For instance, to dereference raw pointers to conduct pointer arithmetic, C’s pointers are converted Rust’s *const or *mut and encapsulated unsafe blocks, thereby reintroducing potential memory errors. Similarly, global variables in C become static mut in Rust, accessed through unsafe blocks, risking race conditions or undefined behavior in multi-threaded contexts. Unsafe operations like transmute further bypass Rust’s type system, leading to potential alignment errors or undefined behavior. Additionally, C2Rust-transpiled code often retains low-level constructs like malloc and free instead of leveraging idiomatic Rust abstractions like Rc, Arc, or Box, increasing reliance on unsafe and requiring developers to manually ensure memory safety.
2.3 Programming language translation with Large Language Models (LLMs)
LLMs offer a new paradigm for programming language translation due to their ability to understand and generate code with contextual awareness while reflecting on their training corpus to adhere to recommended safety coding standards [19, 29, 13]. These features enable LLMs to effectively translate programs between languages. Preliminary studies show that while LLMs perform well on crafted benchmarks such as CodeNet [20], Avatar [1], and EvalPlus [17], they struggle to translate complete software repositories, often failing to capture the complexity and inter-dependencies within real-world projects [13]. This limitation arises from challenges such as ensuring syntactic correctness, maintaining semantic equivalence, and adapting to language-specific features. These shortcomings can lead to issues like incorrect API mappings, type mismatches, and functional errors, underscoring the need for enhanced techniques to improve the reliability and accuracy of LLM-based translations [19].
3 Methodology
In this section, we describe our methodology for translating C code to their safer Rust counterpart, focusing on preserving the intent of the original codebase, but making it more safe. Our tool is designed to operate on complex, real-world C projects, often comprising thousands of lines of code and tens of interdependent files. The translation process prioritizes functional correctness, ensuring that the resulting Rust code behaves as expected under the test cases provided by the user. This section describes our methodology in detail.
3.1 Overview
C2SaferRust is a neuro-symbolic framework designed to translate C code into safer Rust. We leverage on-demand static analysis and LLMs for translation, and differential testing for validating the translated code. However, given the size of typical C project, directly translating an entire C project to Rust in one step is infeasible. We need to come-up with a strategy to translate and validate smaller pieces of code in a step-by-step manner. Thus, we need a framework that allows partial C and partial (translated) Rust code to coexist, as we need to validate the accuracy of the translation by running this hybrid code. To this end, we adopted the C2Rust framework, which transpiles a given C code repository into an equivalent unsafe Rust repository. While the C2Rust-transpiled code retains most of C’s ‘unsafe‘ properties and original intent, it effectively acts as a Rust wrapper, enabling the unsafe code to execute within a Rust environment. Our translation process begins from this starting point, progressively transforming the unsafe Rust into safer Rust.
This unsafe Rust code is then divided into smaller units, called translation units, using program graphs extracted through symbolic reasoning. Each unit is translated into safer Rust using an LLM, which generates safer code while symbolic reasoning ensures accuracy. Compilation and test cases identify errors, prompting targeted analysis to refine context and guide effective repairs. This iterative approach balances efficiency and accuracy, progressively enhancing translation quality to suit the structure and requirements of each project. An overview of the system is shown in Figure 1. C2SaferRust consists of the following components:
  1. 1.
    Translation Orchestrator (§ 3.2): This component, aided by static analysis, orchestrates translation. It breaks the project into smaller translation units that fit within the LLM’s context. It then selects translation units in the optimal order for translation. Translation units are not independent - when one unit is translated, the side effects have to be propagated to all other affected units. The optimal order for translation is moving up the chain of propagated side effects, i.e., translating a unit before translating the units to which its side effects are propagated. This ensures that while the LLM is translating a unit, it can also assimilate the side effects propagated from other units. We will discuss more about this with concrete examples in § 3.2.
  2. 2.
    Static Analyzer (§ 3.3): We collect key dependency data (e.g., call graph, dataflow analysis) using traditional program analysis techniques. Since our translation process is incremental, the program changes at every translation step, and the static analyzer must be re-run each time.
  3. 3.
    Translation (§ 3.4): The translation prompt is composed of the code component to be translated, augmented with information derived from static analysis. This information describes the surrounding context of the code segment and its dependencies, which guides the LLM in translating the unsafe Rust code to safe Rust.
  4. 4.
    Validator (§ 3.4): This component acts as a regression testing framework, validating the translation of each unit by first compiling the new code, and then running system-level tests provided along with the original C code. The safe Rust components replace their unsafe counterparts progressively, allowing both safe and unsafe Rust to coexist during validation.
  5. 5.
    Repair (§ 3.4): After a unit is translated, if there an error in validation, then the repair component of our system collects feedback from compilation and execution. Then, it prompts the LLM with this feedback and asks it to re-generate the translation of the unit.
For each translation unit, the cycle of Validation and Repair continues until the validation succeeds, or we exceed $N$ attempts (where $N$ is a parameter of our system). If we are unable to translate a unit within $N$ attempts, we leave it as its original unsafe Rust version, thus maintaining overall correctness.
In this manner, we go through the entire unsafe Rust program, processing each translation unit in order. Once we translate (or attempt to translate) a unit, we do not come back to it later. In the remainder of this section, we describe each of the components of our system in detail.
3.2 Translation Orchestrator
3.2.1 Decomposing the program into translation units
Our goal is to translate the entire program into safe Rust; however, we are limited by the size of the LLM context window. Further, even if the input and output fit within the input context and output context, LLMs are more prone to hallucination and translation errors when the input is long. Our solution is to decompose the program into translation units of less than $L$ lines each, and translate these units individually.
If a function has fewer than $L$ lines, we extract the whole function as a translation unit. Otherwise, we divide the function into multiple translation units of fewer than $L$ lines, which are possibly nested. At a high level, we traverse the AST of the function in a DFS post-order traversal, and visit only the sub-trees that are greater than $L$ lines long. In each such sub-tree, we form groups of sibling nodes and extract the corresponding statements in each group as a translation unit. Once a unit is extracted, it is removed from the AST. In this way, as we move up the AST, we ensure that every subtree has fewer than $L$ lines. By induction, at the end of the traversal, the root tree will also have fewer than $L$ lines. For the full algorithm, please refer to the appendix.
Figure 2: An example of the output of our decomposition algorithm as described in § 3.2.1, applied to the cat function from the cat program. Translation Unit 1 is the “root” unit that represents the function as a whole, and all other units are nested within it. The units are annotated with comments showing live-in and live-out information, which is described in § 3.3.
Each function will be decomposed into $n$ translation units. These units are of two kinds:
  • Root translation unit: One out of the $n$ translation units corresponds to the entire function, and contains all the other units within it. It also includes the function signature.
  • Inner translation units: The remaining $n-1$ translation units are nested within each other or within the root units.
An example of this is shown in Figure 2. Note that this decomposition does not include code that is outside of functions, like structure and enum definitions. Translating these into safe Rust is a non-trivial problem, and is out of the scope of our paper.
3.2.2 Arranging translation units in an optimal order
The order in which we convert unsafe Rust translation units to safe Rust makes a difference to the quality of translation. Let us say we are converting a function A that calls an unsafe function B. Then even if we “clean up” the body of A to use safe Rust for the most part, it must still rely on unsafe blocks to call B. Further, if B has arguments or return values that are raw pointers, then A will have to use unsafe code to manipulate these values. However, once B is “cleaned up” to use only safe and safer Rust types, then we can convert A to safe Rust in its entirety. So each function should ideally be processed before its callers.
Figure 3: A portion of the callgraph of cat. Each node denotes a function, and the edges denote calls. The functions are sorted in a reverse topological order. The numbers next to each function denote their positional index in the ordering of translation units. main_0 and cat are associated with a range of indices (4-8 and 16-29) because they have multiple translation units within them.
To ensure this, we build a call-graph of the program and compute its weakly connected components. A weakly connected component (WCC) is a subgraph such that, if all its directed edges are converted to undirected edges, all its nodes have paths between each other. Functions within each WCC need to be ordered such that callee comes before caller. So we do a reverse topological sort of each WCC. If a WCC has a cycle, there is no topological ordering. In this case, we start at an arbitrary node in this WCC and order the nodes according to a depth-first post-order traversal.
This gives us an ordering of functions within each WCC of the call graph. The order of functions between two WCCs doesn’t matter, so we process the WCCs one at a time in an arbitrary order to create the full ordering of all functions in the program.
This gives us an ordering of functions; however each function could contain multiple translation units. As we will discuss in § 3.4, the order of translation units within a function does not matter for our current translation strategy. However, we ensure that all the translation units corresponding to a single function are contiguous in the translation order.
3.3 Static Analyzer
We need to annotate each translation unit with enough context information to enable the LLM to translate the code well. Further, each unit cannot be translated in isolation, because it could have ripple effects on other regions of the code. Our solution is to translate all these affected regions simultaneously.
We obtain this information through static analysis of the program. The information we derive differs for a root translation unit and an inner translation unit.
A root translation unit will contain the function header and signature, along with part of the function body. We cannot translate this in isolation, since its parameters or return type might change in this process (e.g., raw pointers might become borrows). These changes are actually desirable, because we need to modify function signatures to transition away from unsafe Rust idioms like raw pointers.
When a function signature changes, this change will spread to all of its call sites. So we ask the LLM to translate a function together with all of its call sites. We also add other contextual information to help the LLM, such as global variables and use statements. Thus, each slice for a root translation unit consists of:
  1. 1
    The source code of the translation unit
  2. 2
    The function’s call sites
  3. 3
    Global variables used in the function
  4. 4
    Module imports used in the file
On the other hand, for an inner unit, we need to tell the LLM about the variables that are used and created within the unit. We use dataflow analysis to get the live-in and live-out variables for the unit. So each slice consists of:
  1. 1
    The source code of the translation unit
  2. 2
    Its live-in and live-out variables
  3. 3
    Global variables used in the translation unit
  4. 4
    Module imports used in the file
Figure 4 shows slices for a root translation unit and an inner translation unit, composed into an LLM prompt.
Figure 4: An example of translating code units with LLM. For the root translation unit on the left, we generate the translation of the function body plus all of its call sites simultaneously. For the inner translation unit on the right, we translate just the code contained within it. In both cases, we also allow the model to specify any additional imports it needs.
3.4 Translation, Validation and Repair
We now attempt to translate each translation unit from unsafe Rust to safe, idiomatic Rust, using an LLM. We pass each unit to an LLM, along with its context slice as described in § 3.3. We instruct the LLM to convert the translation unit to idiomatic Rust. If the unit is a root translation unit of a function and its context contains call sites, we instruct the LLM to translate the call sites too. We specify a particular format for the LLM response that allows us to easily parse it and separate it into function and call-site translations. If the unit is an inner translation unit, we ask the LLM to ensure that the live-in and live-out variables at the beginning and end of the unit are not affected. This means that translating such a translation unit has no side-effects on any other part of the program. This substantiates our claim in § 3.2.2 that the order in which we translate the units within a function does not matter.
Sometimes, the translated unit needs additional use statements to work; we allow the LLM to specify this too in its response. An example of a prompt and LLM response is shown in Figure 4.
We then swap out the original spans of code and swap in the LLM-generated code. We attempt to compile the code and run the test suite. If we encounter an error, we provide it as feedback to the LLM in the same conversation, and ask it to re-generate the entire code, including the function body, call sites and imports. We repeat this for up to $N$ attempts (where $N$ is a parameter of our technique). If one of these $N$ attempts succeeds, we move on to the next function; if all $N$ attempts fail, we restore the original version of the function.
In this manner, we go through the entire program, processing each translation unit exactly once. Throughout this process, we maintain correctness modulo the test cases. Our final program will still have some unsafe functions, but these can be converted to safe Rust by a human expert.
4 Experimental Setup
4.1 Datasets
We contribute a benchmark dataset of 7 C programs collected from GNU coreutils [^2] https://github.com/coreutils/coreutils. These are the basic file, shell and text manipulation utilities of the GNU operating system, and include common shell commands like ls and cat. We chose coreutils because it is a large, mature C code project, with many constituent programs of varying sizes. Further, there is an ongoing community effort to manually rewrite coreutils in Rust [^3] https://github.com/uutils/coreutils, which means that a) translating coreutils to Rust is an important real-world problem, and b) future researchers using our benchmark can compare the quality of their automatically-generated idiomatic Rust code with human-written Rust code. Finally, each benchmark in coreutils comes with end-to-end system-level test cases, which have been carefully constructed using fuzzing to ensure high coverage[4].
The 7 programs we picked from coreutils are split, pwd, cat, truncate, uniq, tail, and head. Running C2Rust on these programs was not straightforward, because C2Rust does not work directly with Makefiles of complex programs. Coreutils uses a compilation pattern where some object files are bundled into static libraries, and those are then linked to the necessary executables. This compilation pattern makes it difficult for C2Rust to create a compilation database, which is necessary for further analysis.
To make this problem tractable, we create simpler individual Makefiles for each of the 7 programs, including only the .c and .h files that are strictly necessary for each program. We collect the source files for each program iteratively. Starting with the main C file for that program (e.g.,  split.c for split), we collect the necessary header files and attempt to build the program. If the linking step fails due to missing symbols, we collect the necessary source files that define these symbols and attempt to build the program again. We repeat until the executable is built successfully. The simpler Makefile is then fed to C2Rust.
The statistics of these 7 programs are shown in Table I. In addition to these 7 programs from coreutils, we also use 10 benchmark programs from a prior work, Laertes [8]. These do not come with end-to-end test cases. Although some of the utilities have unit tests, this is not useful because we cannot run C unit tests on Rust programs.
Program # LoC # Func Max Fn Len Mean Fn Len
Coreutils split 13848 207 1634 52.95
pwd 5859 127 902 33.02
cat 7460 166 1006 33.84
truncate 7181 124 902 44.65
uniq 8299 167 902 36.26
tail 14423 266 902 43.81
head 8047 153 902 40.77
Laertes tulipindicators 44486 234 306 32.10
xzoom 2524 11 536 134.09
genann 2084 32 201 52.44
optipng 95560 576 42124 143.63
urlparser 1118 22 409 38.41
lil 5400 160 283 30.44
qsort 41 3 17 10.67
snudown 6521 92 792 63.53
grabc 1046 7 61 31.57
bzip2 43374 128 3289 98.84
TABLE I: The statistics of the 7 programs in our coreutils benchmark dataset, plus the 10 programs from Laertes.
4.2 Baselines
We evaluate against two previous approaches for making C2Rust output more natural.
  • Laertes [8]: Laertes uses a set of simple transformation rules to rewrite each raw pointer as a borrow, and then uses compiler feedback to fix errors until the program compiles.
  • CROWN [30]: CROWN derives ownership constraints for each raw pointer based on their usage patterns, and uses an SMT solver to solve these constraints and transform each raw pointer into a borrow.
4.3 Implementation
We implement C2SaferRust on the nightly-2022-08-08 version of Rust. It is implemented as a cargo subcommand, so it can be run on full Cargo packages. We use the Rust compiler’s High-level Intermediate Representation (HIR) to construct the program call graph and to decompose the program into chunks. The AST traversal in Algorithm 1 is implemented as an instance of the rustc_hir::intravisit::Visitor trait. The slicing described in § 3.3 is also performed on the HIR.
We use the Mid-level Intermediate Representation (MIR) to solve the dataflow analysis for each chunk as described in § 3.3. We implement our analysis as an instance of the GenKillAnalysis trait from rustc_mir_dataflow, which computes live-in and live-out variables for a basic block. We use the MIR debug information to find the first and last basic blocks in a chunk, and get the live-in variables of the first block and the live-out variables of the last block. These become the live-in and live-out variables of the chunk.
For all our experiments, the LLM used is gpt4o-mini by OpenAI. We fixed the maximum chunk size, $L$, at 150 lines, and the number of LLM feedback iterations, $N$, at 5. In this setting, the smallest baseline took under 5 mins to translate, while the largest took nearly 14 hours.
4.4 Evaluation
We use two metrics from prior work to evaluate the quality of translation:
  • Raw Pointer Declarations: The number of raw pointer declarations, *const T or *mut T, in the entire program. This includes declarations in function signatures and within structures. This metric was used to evaluate Laertes [8].
  • Raw Pointer Dereferences: The number of raw pointer dereferences throughout the program. This includes all dereferences of the form *<expr> where <expr> evaluates to a raw pointer type (verified by type checking). This metric was also used to evaluate Laertes [8].
Raw pointers are not a feature of idiomatic Rust, and the fewer raw pointers in a program, the more natural it is. Further, dereferencing raw pointers is an unsafe operation which could lead to segmentation faults and other undefined behavior if not used correctly. So for both of the above metrics, a lower value indicates safer code. We further analyze unsafe blocks by investigating their length and types of statements, through the following three metrics:
  • Unsafe Lines of Code: The total lines of code (LOC) contained within all unsafe blocks in the program. This includes the bodies of unsafe functions, which are also treated as unsafe blocks for this metric. A lower value for unsafe LOC indicates a more natural and idiomatic Rust program, as it minimizes the reliance on unsafe constructs.
  • Unsafe Type Casts: Type casts, within unsafe blocks, especially those involving raw pointers (such as ptr as *const u8) are a common pattern which bypass Rust’s strict type system, allowing for potential misuse and may result in undefined behavior if the cast is used incorrectly.
  • Unsafe Call Expressions: Function calls within unsafe blocks, such as std::alloc::alloc(layout) or other custom unsafe functions, allow for a direct interaction with low-level operations. This can be dangerous when encapsulated in unsafe blocks because they bypass Rust’s safety guarantees and potentially lead to memory corruption, dangling pointers, or other undefined behavior.
5 Results
5.1 Effectiveness of C2SaferRust in reducing unsafe constructs (RQ1)
Motivation: In the first research question, we try to evaluate how effectively C2SaferRust can transform C2Rust’s unsafe code into safe and idiomatic Rust.
Evaluation: We run C2SaferRust on each program in our benchmark dataset, which was compiled from coreutils. For each program, we calculated the following metrics both before and after translation (see § 4.4 for detailed information): (1) the count of raw pointer declarations, (2) the number of raw pointer dereferences, and (3) the extent of unsafe lines of code. Additionally, our analysis was expanded to encompass measurements of (1) unsafe type casts, (2) unsafe call expressions, and (3) unsafe let declarations.
TABLE II: Translating CoreUtils to safer Rust using C2SaferRust. Overall results suggest C2SaferRust offers enhanced safety by reducing reliance on raw pointers and unsafe operations.
[…]
Before
After
$\Delta\%$
Before
After
[…]
(a) Number of raw pointer declarations and dereferences across all applications before and after translation.
[…]
Before
After
$\Delta\%$
Before
After
$\Delta\%$
Before
After
[…]
(b) Number of unsafe expressions and declarations AST nodes across CoreUtils projects before and after translation.
Results: § 5.1 tabulates the effectiveness of translating CoreUtils projects into safer Rust using C2SaferRust. In the following, we discuss the key findings in three key metrics: raw pointer declarations, raw pointer dereferences, and line of unsafe code.
  1. 1.
    Raw Pointer Declarations. We observe an average reduction of raw pointer declarations by $\mathbf{25\%}$ with notable reductions across all projects. The most significant improvement is observed in cat, which shows a $\mathbf{38\%}$ reduction in raw pointer declarations.
  2. 2.
    Raw Pointer Dereferences. Unsafe raw pointer dereferences are reduced on average by $\mathbf{22\%}$, demonstrating the tool’s effectiveness in minimizing unsafe manipulations. The project uniq shows the largest decrease at $\mathbf{27\%}$, indicating a significant impact on pointer-heavy codebases.
We also observed a notable reduction in unsafe code (§ 5.1) across all CoreUtils projects after translation to safer Rust.
  1. 1.
    Lines of unsafe code. Lines of unsafe code decrease by an average of $\mathbf{24\%}$ across all projects, reflecting a significant enhancement in code safety and idiomaticity. uniq shows the largest reduction of $\mathbf{28\%}$, highlighting C2SaferRust’s ability to minimize unsafe code.
  2. 2.
    Unsafe Type Cast: Unsafe type casts decrease, but by a relatively smaller amount compared to the other metrics - an average of $\mathbf{8\%}$ reduction. The most substantial decrease is in head, from 1378 to 1189 casts (by $\approx\mathbf{14}\%$). The sub-par performance on this metric can be attributed to the extra casts needed to convert raw pointers to other types and back, e.g., 2 and 3 in Fig. 5. This is discussed in more detail below.
  3. 3.
    Unsafe Call Expression: Unsafe function calls in projects decrease by an average of 24%. The most notable reduction in calls is in uniq, from 6066 to 4397 (by $\approx\mathbf{28}\%$), highlighting the tool’s capacity to refactor and mitigate risks from low-level APIs.
(a) Original C implementation
(b) C2SaferRust output
(c) C2Rust translation
Figure 5: C code versus C2Rust code versus LLM-translated safer Rust, for the simple_cat function from the cat program.
Discussion: Figure 5 illustrates a function using the original C code (Fig. 5), C2Rust translation (Fig. 5), and C2SaferRust’s safer output (Fig. 5) for comparison. We make the following observations:
  1. 1)
    C2SaferRust uses idiomatic and native Rust types. Non-native types in the C2Rust code (such as libc::c_char, see 1 in Fig. 5) are replaced with their native Rust types (i.e., &mut [u8] as seen in 1, Fig. 5). Note that in some cases these pointers are cast back into their raw pointer-type when passed as an argument to other functions, e.g., 2 and 3 in Fig. 5.
  2. 2)
    C2SaferRust continues to make use of FFI calls. C2SaferRust accomplishes certain core functionalities through calls to C API functions via foreign function interfaces (FFIs), e.g., safe_read and error (see 2 and 3 in Fig. 5). This limits the extent of code safety since the arguments and return types of these functions still use raw pointer types. These may only be eliminated when all the external functions are replaced with safe Rust equivalents; this is beyond the scope of this work.
  3. 3)
    C2SaferRust eliminates nonstandard control flow structures and dead code. Rust code generated by the C2Rust transpiler generates a large number of unreachable branches in the C2Rust code. It makes explicit calls to branch conditions that cannot be executed (e.g., if 0 != 0 as seen in 4, Fig. 5). C2Rust also uses certain macros such as unreachable!() (as seen in 5, Fig. 5). C2SaferRust effectively eliminates these dead code fragments and produces more concise and idiomatic Rust code.
{myshadowbox}
Summary: C2SaferRust reduces unsafe constructs in CoreUtils, cutting raw pointers by 38% and dereferences by 27%. Unsafe code lines, type casts and call expressions decrease by 24%, 8% and 24%, enhancing Rust safety and idiomaticity.
(b) #Attempts Required to Translate a Function, vs Length
Figure 6: For our coreutils benchmark dataset - a plot of the fraction of successful translations versus function length, and the number of attempts required to translate a function versus function length. This includes only functions that are translated as a single unit, i.e., functions that are under 150 lines.
Results: The results for the component ablation are shown in Table IV. There are 3 utilities and 5 metrics, which makes a total of 15 settings. C2SaferRust outperforms both “No Chunking” as well as “Random Order” in 9 out of 15 settings, showing a clear benefit of both design choices.
LABEL:fig:success_vs_length shows the rate of translation success versus function length, and Figure 5(b) shows the number of attempts needed to translate a function vs its length. If we ignore the outliers to the right of the distribution, we can see that as the function length increases, the translation success rate decreases and the number of required attempts increases. This clearly indicates that the longer a function is, the harder it is for an LLM to translate it. This provides empirical support for our decision to break up very long functions into smaller chunks for translation.
Program Method Raw Ptr Raw Ptr Unsafe Unsafe Unsafe
Decls $\downarrow$ Derefs $\downarrow$ Lines $\downarrow$ Calls $\downarrow$ Casts $\downarrow$
cat Original 192 317 5625 1038 5625
No Chunking 137 248 3638 922 3638
Random Order 139 246 4194 955 4194
Our Tool 120 240 4189 912 4189
uniq Original 227 343 6066 1150 6066
No Chunking 169 252 4382 1114 4382
Random Order 164 258 4516 1117 4516
Our Tool 166 250 4397 1025 4397
head Original 192 442 6245 1378 6245
No Chunking 142 360 4814 1380 4814
Random Order 142 349 4735 1252 4735
Our Tool 141 351 4699 1189 4699
TABLE IV: Showing the contribution of the components of our method by ablation on three selected benchmarks.
6 Related Work
Converting C to unsafe Rust. Corrode [6] and C2Rust [10] are two systems that aim to convert C code to functionally equivalent Rust code. Both tools have limitations and corner cases that they cannot handle. The primary difference between them is that C2Rust is still being actively developed and maintained, while Corrode has not been updated for several years. The Rust code that these tools produce is non-idiomatic and makes extensive use of unsafe features. Citrus [5] is a third tool that uses syntactic mappings to convert C to non-idiomatic, unsafe Rust. However, unlike Corrode and C2Rust, Citrus doesn’t aim to produce functionally equivalent code. The translated code may not run or compile. Citrus aims to be a lighter alternative to C2Rust, to produce a first draft of Rust code that can manually be refactored into a working Rust program.
Converting unsafe Rust to safer Rust. References in Rust are tracked by the borrow checker to ensure that they don’t extend beyond the scope of the value they are pointing to. However, the non-idiomatic Rust code generated by C2Rust makes extensive use of raw pointers, which are not tracked by the borrow checker. There has been some work on automatically converting a subset of the raw pointers in a program to references, thus creating safer Rust code.
Laertes [8] was the first such approach, which used compiler errors to extract information about lifetimes and mutability of each raw pointer, and tried to iteratively rewrite each one as a borrow. The authors also performed an in-depth study about how the different ways in which unsafe features are used in C2Rust transpiled code. However, this method can only be applied to functions where the cause of unsafety is due to a lack of ownership information on a raw pointer. Only 20% of the functions fall into this category, and they are only able to convert about 8.3% of the raw pointers in their benchmark dataset to references. In a follow-up paper [9], the authors of Laertes try to expand their technique to include more categories of raw pointers. With this new approach, they are able to convert about 21% of the pointers to references.
CROWN [30] is another approach that starts with C2Rust code and tries to convert each raw pointer into a reference. It derives constraints for each raw pointer, and solves these constraints using an SMT solver to convert each raw pointer into a reference. However, like Laertus, CROWN too has a limitation - it can handle only mutable, non-array raw pointers. It succeeds at converting about 37% of these to references.
Hong and Ryu [11] develop a tool to perform a very specific type of transformation on C2Rust code - replacing output parameters that indicate success/failure with canonical Rust types like Option and Result. In their subsequent work [12], they study another specific type of transformation on C2Rust code - replacing C-style unions with Rust tagged unions.
LLMs for Code Translation. Roziere et al. [21] proposed TransCoder, an early attempt to tackle code translation using transformer models. They proposed Back Translation, a fully unsupervised approach. Ahmad et al. [2] refined Back Translation with the addition of a code summarization task. In contrast to unsupervised learning, supervised learning requires aligned parallel data across pairs of programming languages, which is challenging to obtain at scale. Roziere et al. [22] addressed this challenge by generating synthetic aligned data using another language model and filtering out incorrect translations using generated unit tests. The last few years have seen the development of LLMs for code and natural language, that are pretrained on massive amounts of data with an autoregressive objective function. Pan et al. [19] and Yang et al. [29] perform thorough empirical evaluations of the translation ability of these models. Yang et al. [29] also propose UniTrans, a tool that generates test cases, provides them along with the code to an LLM for translation, and uses test feedback to iteratively repair the generated translation. Nitin et al. [18] propose to generate multi-modal specifications with a self-consistency filter and provide these along with the code to an LLM for translation. [28] propose to verify candidate LLM translations against the output of a transpiler using rigorous equivalence proving.
However, all of these approaches deal only with settings where the entire program is relatively short and fits within a single LLM context window. The typical setting is that of solutions to competitive programming problems. Shiraishi and Shinagawa [25] was the first work to attempt translation of arbitrarily large programs using LLMs. They slice a C repository into smaller translation units, translate each unit to Rust using an LLM, and use compiler feedback to fix translations. However, they do not use test feedback, and they require each translation unit to be independently compilable.
There have been a few prior works that attempt to repository-level translation for other pairs of languages. Ibrahimzada et al. [13] attempt translation of full Java projects to Python. Zhang et al. [31] translate Go to Rust, focusing on avoiding runtime errors in the translated code. While preparing this manuscript, we came across a recent ArXiv preprint [24] that presents an approach to perform full repository C to Rust translation. They divide the program into translation units, translate each unit using an LLM, and use dynamic analysis along with LLM-based test generation to verify the correctness of each translation. However, our approach is different in certain important aspects. In particular, a) we use a transpiler to provide a skeleton for the target language code, thus allowing us to check end-to-end correctness at every step; b) we break down code up to a sub-method level, ensuring that we can scale to programs with methods of arbitrary length.
7 Threats to Validity
Our metrics are intended to model the safety and naturalness of the code; however it is possible that an improvement in a metric doesn’t correlate with improved safety or naturalness of the code. For example, the number of lines of code within unsafe blocks could decrease simply by eliminating line breaks and relying on semicolons to delimit lines. We therefore assume that the tools we consider do not attempt to game the metrics by performing such manipulations.
The code we generate still uses a significant amount of unsafe code. Some of this is because we do not translate code outside functions, like structure definitions, and interacting with these structures could necessitate the use of some unsafe code. Another observation is that the LLM does not replace C FFI function calls with their safe Rust equivalents; rather, it uses unsafe code to wrap these FFI calls.
8 Conclusion
In this work, we have presented C2SaferRust, a system to automatically translate full C repositories to safer Rust. Based on the novel insight that the C2Rust transpiler can be used as an intermediary to bridge between C and Rust for the purpose of LLM-aided translation, we first convert the entire repository into unsafe Rust. We propose an algorithm to divide the code repository into smaller “translation units”, and extract each translation unit along with its context information. We use an LLM to translate each unit to safer Rust, swap in this translation, and propagate the changes to the rest of the code base.
We contribute a dataset of 7 C programs from GNU CoreUtils with their corresponding C2Rust transpiled versions. Each program comes with end-to-end test cases, allowing one to empirically verify the correctness of translation. On this dataset, we show that we reduce raw pointer declarations and dereferences by up to 38%, and reduce the amount of unsafe code by up to 28%. We also compare our tool with two baselines for making Rust code safer, and observe consistent gains. Future work could look at translating code outside functions, like structure definitions, as well as remapping C FFI calls to safe Rust equivalents.
Appendix A Decomposition Algorithm
1input: G - the AST of the program
2input: L - the max length of a translation unit
3
4output: units - a list of pieces of code, each
5 having $\leq$ L lines
6
7units := []
8walk_subtree(G)
9
10fn walk_subtree(node):
11 for child in node.children:
12 match child.type:
13 Func => visit_func(child),
14 Block => visit_block(child),
15 _ => walk_subtree(child)
16
17fn visit_func(func):
18 // Traverse each function longer than L lines
19 if line_count(func) <= L:
20 chunks.push(func)
21 else:
22 walk_subtree(func)
23
24// A block is any entity within {..} parantheses
25fn visit_block(block):
26 // Walk each subtree longer than L lines
27 for stmt in block.statements:
28 if line_count(stmt) > L:
29 walk_subtree(stmt)
30 lower, upper := 0, 0 // Maintain two cursors
31 while upper $<$ block.statements.len():
32 if line_count(block.statements[lower..upper]) > L:
33 // Adding this statement caused the
34 // length to spill over
35 unit := block.statements[lower..upper-1]
36 units.push(unit)
37 // Remove this unit from the AST
38 G.prune_nodes(unit)
39 lower = upper
40 else:
41 upper += 1
42 // We have some leftover lines
43 if lower $<$ block.statements.len():
44 unit := block.statements[lower..]
45 units.push(unit)
46 // Remove this unit from the AST
47 G.prune_nodes(unit)
Algorithm 1 Decomposing the program into translation units, each having fewer than $L$ lines.