### v2

# C2SaferRust: Transforming C Projects into Safer Rust with NeuroSymbolic Techniques

Abstract
In recent years, there has been a lot of interest in converting C code to Rust, to benefit from the memory and thread safety guarantees of Rust. C2Rust is a rule-based system that can automatically convert C code to functionally identical Rust, but the Rust code that it produces is non-idiomatic, i.e., makes extensive use of unsafe Rust, a subset of the language that doesn’t have memory or thread safety guarantees. At the other end of the spectrum are LLMs, which produce idiomatic Rust code, but these have the potential to make mistakes and are constrained in the length of code they can process. In this paper, we present C2SaferRust, a novel approach to translate C to Rust that combines the strengths of C2Rust and LLMs. We first use C2Rust to convert C code to non-idiomatic, unsafe Rust. We then decompose the unsafe Rust code into slices that can be individually translated to safer Rust by an LLM. After processing each slice, we run end-to-end test cases to verify that the code still functions as expected. We also contribute a benchmark of 7 real-world programs, translated from C to unsafe Rust using C2Rust. Each of these programs also comes with end-to-end test cases. On this benchmark, we are able to reduce the number of raw pointers by up to 38%, and reduce the amount of unsafe code by up to 28%, indicating an increase in safety. The resulting programs still pass all test cases. C2SaferRust also shows convincing gains in performance against two previous techniques for making Rust code safer.
1 Introduction
The Rust language has emerged as a popular alternative to C, thanks to its strong memory safety and thread safety guarantees. Rust uses a memory model based on ownership and borrowing, which prevents accesses of a memory location after it has been de-allocated or freed. It also disallows mutable aliasing, which prevents data races in multi-threaded programs. At the same time, Rust has comparable runtime performance to C, making it suitable for performance-critical applications like device drivers and embedded systems.
In contrast, C code can contain both memory safety and thread safety errors, which makes it prone to security vulnerabilities. For all of these reasons, there has been a lot of interest in converting all C code to Rust. However, doing this manually is challenging and requires tremendous manpower and expertise. Just a single project, the Linux kernel, contains more than 7 million lines of C code, spread over more than 16,000 files. Thus, there is a need to develop tools to automatically convert C to Rust.
Limitations of Existing Approaches: Existing approaches for code translation fall into two broad categories. The first is transpilers or cross-compilers, which are rule-based systems. For instance, the C2Rust project [9] is a transpiler that converts C code to Rust. Although the Rust code produced by C2Rust is almost always correct, it is unsafe Rust, a subset of the language that lacks the memory and thread safety guarantees usually associated with Rust. Further, this code is non-idiomatic, i.e.,, very different from the kind of Rust code usually written by humans, making it hard to read and maintain.
There have been a few approaches [7, 30, 10, 11] that augment the transpiler framework with specific rule-based transformations, such as converting certain datatypes to other types. However, these are limited in scope, and cannot “zoom out” to transform the code at a higher granularity.
The second category of approaches is using Large Language Models (LLMs). These models are trained on vast amounts of human-written code, and thus produce code that is idiomatic, i.e., uses the most conventional idioms and constructs of a programming language. However, this code frequently has errors and is unreliable. This problem is compounded because LLM-generated code has the appearance of correctness, which makes any errors harder to spot.
Another challenge when applying LLMs for code translation is the length of input and/or output code. LLMs operate on a fixed context window, which means that they can only process inputs and generate outputs that are less than a certain predefined length. A real-world code project with dozens of files and hundreds of lines per file may not fit within this window. Further, even if the entire project does fit, it is known that LLMs tend to produce more errors as the length of the input or output increases [18]. For these reasons, most LLM-based code translation research has focused on programs that are relatively short standalone files, such as those used in competitive programming contests.
Our Approach: In this paper, we combine the strengths of the transpiler approach and the LLM approach to design C2SaferRust- an algorithm to convert full real-world C projects into Rust. Tackling an entire codebase at once can lead to inefficiencies and increased complexity. Instead, we propose breaking down the translation process into smaller, manageable components, such as individual methods or functions. These smaller units can be translated first, and once translated, the changes can be progressively integrated into the larger project by carefully propagating the altered parameters, variables, and interfaces to the rest of the codebase. This requires strategic planning to ensure that dependencies between components are maintained and that changes flow seamlessly throughout the system.
C2SaferRust uses the C2Rust transpiler in the first step to convert all the C code to non-idiomatic unsafe Rust. Then, we incrementally convert slices of this code to safe Rust using an LLM. The transpiled unsafe Rust code thus acts as a bridge between the original C code and the LLM-generated safe Rust code. We traverse the parse tree of the transpiled code to decompose it into smaller units which can be individually translated.
We use static program analysis to analyze dependencies between these units and arrange them in an optimal order for translation. Then we augment each unit with context information, again derived using static analysis, and prompt an LLM with this information to translate the code into safe Rust. We ensure that each unit is translated correctly by compiling the code and running end-to-end test cases each time an LLM produces a translation. If there are errors, we provide the compiler or test feedback to the LLM and iteratively refine the translations.
Figure 1: An overview of our system, C2SaferRust, to translate a C program to safer Rust. We first run the C2Rust transpiler to get non-idiomatic Rust, then decompose the program into chunks. For each chunk, we slice the program to extract relevant context information and dataflow information, and compose this into a translation prompt for an LLM. The LLM produces safer Rust, which we swap back in and test end-to-end. If the program doesn’t compile or pass test cases, we give the LLM feedback and repeat.
Summary of Results: We evaluate on two datasets - one contributed by us, and one from a prior work, Laertes [7].
  • On our dataset, C2SaferRust is able to reduce the amount of unsafe code by up to 28%. We are also able to substantially reduce the occurrence of raw pointers, a datatype that is exclusively associated with unsafe Rust.
  • On the Laertes dataset, C2SaferRust outperforms two selected baseline approaches on a majority of the programs for all of our metrics.
  • We show through an ablation study that each of the components of our design contributes to its overall performance.
Our Contributions: In summary, our contributions in this paper as as follows:
  1. 1.
    We adopt a novel approach of using C2Rust transpiled code as a bridge between C code and LLM-generated safe Rust, combining the strengths of rule-based transpiler approaches and LLMs. We traverse the program parse tree to decompose a large program into smaller units that can be individually translated by an LLM. We then augment each unit with context information derived from static program analysis. We test and validate the LLM-generated translations, and use feedback-guided repair to iteratively refine them.
  2. 2.
    We contribute a benchmark dataset of 7 programs consisting of utilities from the GNU Coreutils project. Each program comes with end-to-end system-level test cases, enabling us to evaluate correctness.
  3. 3.
    The ecosystem of static analysis tools for Rust is rather limited, and so we had to implement most of our analyses as Rust compiler passes, directly using the Rust compiler’s intermediate representations. This lays the groundwork to build future Rust static analysis tools. For instance, we build program slicing and callgraph analysis algorithms, which are quintessential static analysis techniques. We make all of this code publicly available on GitHub [^1] https://github.com/vikramnitin9/c2saferrust.
2 Background
2.1 Safety and Unsafety in Rust
Rust is a systems programming language designed to offer memory safety guarantees while providing performance comparable to C and C++. In contrast, C and C++ are widely used but prone to memory errors and security vulnerabilities [6, 13, 23]. Rust achieves safety through a strict ownership model: each value in memory has a single owner, and deallocation occurs automatically when the owner goes out of scope. Values may be borrowed by other variables, but such borrows are annotated with lifetimes, which the compiler tracks to ensure validity. A common pattern is for functions to borrow values from callers for the duration of the function and return ownership afterward.
Rust’s type system distinguishes between immutable (&T) and mutable (&mut T) borrows, enforcing exclusivity rules to prevent data races. Multiple immutable borrows may coexist, as they do not modify state, but a mutable borrow requires exclusive access, disallowing any simultaneous borrows. The borrow checker ensures statically that these constraints hold, resulting in memory safety without runtime overhead [27].
While these guarantees eliminate entire classes of bugs, they also make certain programming tasks more difficult. In particular, designing data structures that rely on shared ownership or cyclic references (such as graphs) becomes non-trivial, and careful design is needed to avoid lifetime conflicts. Safety enforcement may also introduce bounds checks or borrow checking overhead and complicate interactions with low-level system components or foreign libraries [3]. To support such cases, Rust exposes the unsafe keyword, permitting raw pointer manipulation, mutable static memory, and calls to unsafe functions (including FFI and compiler intrinsics) [26]. Unsafe enables performance-critical or low-level operations but transfers responsibility for correctness from the compiler to the programmer.
The option to selectively use unsafe while retaining strong guarantees elsewhere has driven efforts to port legacy C and C++ systems into Rust, aiming to reduce vulnerability surfaces associated with manual memory management [23]. However, this process is often labor-intensive. The permissive memory model in C—with raw pointers, unchecked array access, and no concept of ownership—makes straightforward translation into Rust’s stricter semantics challenging and error-prone [14].
2.2 Translating C to Rust using Transpilation
A common approach to porting C code is the use of source-to-source compilers, or transpilers, which aim to transform code from one language into another while preserving semantics. Tools such as C2Rust [9] translate C constructs into Rust systematically and can automatically produce compilable Rust code.
Challenges.  Despite providing a starting point, transpilation often produces code that compiles but does not fully benefit from Rust’s safety guarantees. Transpilers map raw pointers to *const and *mut, frequently requiring unsafe dereferencing, and global variables in C typically become static mut, which also demands unsafe access. Common C idioms such as pointer arithmetic, manual memory management, or type reinterpretation via transmute reintroduce the possibility of undefined behavior. Moreover, transpiled code often retains low-level malloc/free semantics rather than adopting idiomatic Rust abstractions like Box, Rc, or Arc, increasing reliance on unsafe and complicating verification of memory safety.
2.3 Programming language translation with Large Language Models (LLMs)
Large Language Models offer an emerging alternative for code translation. Their ability to generate code with contextual reasoning allows them to restructure programs more flexibly than rule-based transpilers, and they may incorporate safe coding idioms learned from training corpora [17, 29, 12]. LLMs have shown promising results on structured benchmarks such as CodeNet [20], Avatar [1], and EvalPlus [15], where problems are self-contained and relatively small.
However, studies demonstrate that translating real-world codebases remains difficult. LLMs often struggle to maintain semantic equivalence across multi-file projects, preserve API interactions, or reason about complex type and memory relationships [12]. These limitations manifest as type mismatches, missing edge cases, broken dependencies, or unsafe substitutions. As a result, improving robustness and safety in LLM-based translation remains an active research direction [17].
3 Methodology
3.1 Overview
C2SaferRust is a neuro-symbolic framework for translating C programs to their safer Rust counterparts, focusing on preserving the intent of the original code. C2SaferRust is designed to operate on complex, real-world C projects, often comprising thousands of lines of code and tens of interdependent files. The translation process prioritizes functional correctness, ensuring that the resulting Rust code behaves as expected under the test cases provided by the user.
We leverage on-demand static analysis and LLMs for translation, and differential testing for validating the translated code. However, given the size of typical C project, directly translating an entire C project to Rust in one step is infeasible. We need to come-up with a strategy to translate and validate smaller pieces of code in a step-by-step manner. Thus, we need a framework that allows partial C and partial (translated) Rust code to coexist, as we need to validate the accuracy of the translation by running this hybrid code. To this end, we adopted the C2Rust framework, which transpiles a given C code repository into an equivalent unsafe Rust repository. While the C2Rust-transpiled code retains most of C’s ‘unsafe‘ properties and original intent, it effectively acts as a Rust wrapper, enabling the unsafe code to execute within a Rust environment. Our translation process begins from this starting point, progressively transforming the unsafe Rust into safer Rust.
This unsafe Rust code is then divided into smaller units, called translation units, using program graphs extracted through symbolic reasoning. Each unit is translated into safer Rust using an LLM, which generates safer code while symbolic reasoning ensures accuracy. Compilation and test cases identify errors, prompting targeted analysis to refine context and guide effective repairs. This iterative approach balances efficiency and accuracy, progressively enhancing translation quality to suit the structure and requirements of each project. An overview of the system is shown in Figure˜1. C2SaferRust consists of the following components:
  1. 1.
    Translation Orchestrator (§˜3.2): This component, aided by static analysis, orchestrates translation. It breaks the project into smaller translation units that fit within the LLM’s context. It then selects translation units in the optimal order for translation. Translation units are not independent - when one unit is translated, the side effects have to be propagated to all other affected units. The optimal order for translation is moving up the chain of propagated side effects, i.e., translating a unit before translating the units to which its side effects are propagated. This ensures that while the LLM is translating a unit, it can also assimilate the side effects propagated from other units. We will discuss more about this with concrete examples in §˜3.2.
  2. 2.
    Static Analyzer (§˜3.3): We collect key dependency data (e.g., call graph, dataflow analysis) using traditional program analysis techniques. Since our translation process is incremental, the program changes at every translation step, and the static analyzer must be re-run each time.
  3. 3.
    Translation (§˜3.4): The translation prompt is composed of the code component to be translated, augmented with information derived from static analysis. This information describes the surrounding context of the code segment and its dependencies, which guides the LLM in translating the unsafe Rust code to safe Rust.
  4. 4.
    Validator (§˜3.4): This component acts as a regression testing framework, validating the translation of each unit by first compiling the new code, and then running system-level tests provided with the original C code. The safe Rust components replace their unsafe counterparts progressively, allowing safe and unsafe Rust to coexist during validation.
  5. 5.
    Repair (§˜3.4): After a unit is translated, if there an error in validation, then the repair component of our system collects feedback from compilation and execution. Then, it prompts the LLM with this feedback and asks it to re-generate the translation of the unit.
For each translation unit, the cycle of Validation and Repair continues until the validation succeeds, or we exceed $N$ attempts (where $N$ is a parameter of our system). If we are unable to translate a unit within $N$ attempts, we leave it as its original unsafe Rust version, thus maintaining overall correctness. In this manner, we go through the entire unsafe Rust program, processing each translation unit in order. Once we translate (or attempt to translate) a unit, we do not come back to it later. In the remainder of this section, we describe each of the components of our system in detail.
3.2 Translation Orchestrator
3.2.1 Decomposing the program into translation units
Our goal is to translate the entire program into safe Rust; however, we are limited by the size of the LLM context window. Further, even if the input and output fit within the context window, LLMs are more prone to hallucinations and translation errors when the input is long. Our solution is to decompose the program into translation units of less than $L$ lines each, and translate these units individually.
If a function has fewer than $L$ lines, we extract the whole function as a translation unit. Otherwise, we divide the function into multiple translation units of fewer than $L$ lines, which are possibly nested. At a high level, we traverse the AST of the function in a DFS post-order traversal, and visit only the sub-trees that are greater than $L$ lines long. In each such sub-tree, we form groups of sibling nodes and extract the corresponding statements in each group as a translation unit. Once a unit is extracted, it is removed from the AST. In this way, as we move up the AST, we ensure that every subtree has fewer than $L$ lines. By induction, at the end of the traversal, the root tree will also have fewer than $L$ lines. For the full algorithm, please refer to the appendix.
Figure 2: An example of the output of our decomposition algorithm as described in §˜3.2.1, applied to the cat function from the cat program. Translation Unit 1 is the “root” unit that represents the function as a whole, and all other units are nested within it. The units are annotated with comments showing live-in and live-out information, which is described in §˜3.3.
Each function will be decomposed into $n$ translation units. These units are of two kinds:
  • Root translation unit: One out of the $n$ translation units corresponds to the entire function, and contains all the other units within it. It also includes the function signature.
  • Inner translation units: The remaining n1n-1 translation units are nested within each other or within the root units.
An example of this is shown in Figure˜2. Note that this decomposition does not include code that is outside of functions, like structure and enum definitions. Translating these into safe Rust is a non-trivial problem, and is out of the scope of our paper.
3.2.2 Arranging translation units in an optimal order
The order in which we convert unsafe Rust translation units to safe Rust makes a difference to the quality of translation. Let us say we are converting a function A that calls an unsafe function B. Then even if we “clean up” the body of A to use safe Rust for the most part, it must still rely on unsafe blocks to call B. Further, if B has arguments or return values that are raw pointers, then A will have to use unsafe code to manipulate these values. However, once B is “cleaned up” to use only safe and safer Rust types, then we can convert A to safe Rust in its entirety. So each function should ideally be processed before its callers.
Figure 3: A portion of the callgraph of cat. Each node denotes a function, and the edges denote calls. The functions are sorted in a reverse topological order. The numbers next to each function denote their positional index in the ordering of translation units. main_0 and cat are associated with a range of indices (4-8 and 16-29) because they have multiple translation units within them.
To ensure this, we build a call-graph of the program and compute its weakly connected components. A weakly connected component (WCC) is a subgraph such that, if all its directed edges are converted to undirected edges, all its nodes have paths between each other. Functions within each WCC need to be ordered such that callee comes before caller. So we do a reverse topological sort of each WCC. If a WCC has a cycle, there is no topological ordering. In this case, we start at an arbitrary node in this WCC and order the nodes according to a depth-first post-order traversal.
This gives us an ordering of functions within each WCC of the call graph. The order of functions between two WCCs doesn’t matter, so we process the WCCs one at a time in an arbitrary order to create the full ordering of all functions in the program. Further, each function could contain multiple translation units. As we will discuss in §˜3.4, the order of translation units within a function does not matter for our current translation strategy. However, we ensure that all the translation units corresponding to a single function are contiguous in the translation order.
3.3 Static Analyzer
We need to annotate each translation unit with enough context to enable the LLM to translate the code well. Further, each unit cannot be translated in isolation, because it could have ripple effects on other regions of the code. Our solution is to translate all these affected regions simultaneously.
We obtain this information through static analysis of the program. The information we derive differs for a root translation unit and an inner translation unit.
A root translation unit will contain the function header and signature, along with part of the function body. We cannot translate this in isolation, since its parameters or return type might change in this process (e.g., raw pointers might become borrows). These changes are actually desirable, because we need to modify function signatures to transition away from unsafe Rust idioms like raw pointers.
When a function signature changes, this change will spread to all of its call sites. So we ask the LLM to translate a function together with all of its call sites. We also add other contextual information to help the LLM, such as global variables and use statements. Thus, each slice for a root translation unit consists of:
  1. 1
    The source code of the translation unit
  2. 2
    The function’s call sites
  3. 3
    Global variables used in the function
  4. 4
    Module imports used in the file
On the other hand, for an inner unit, we need to tell the LLM about the variables that are used and created within the unit. We use dataflow analysis to get the live-in and live-out variables for the unit. So each slice consists of:
  1. 1
    The source code of the translation unit
  2. 2
    Its live-in and live-out variables
  3. 3
    Global variables used in the translation unit
  4. 4
    Module imports used in the file
Figure˜4 shows slices for a root translation unit and an inner translation unit, composed into an LLM prompt.
Figure 4: Translating units with an LLM. For the root translation unit (left), we generate the translation of the function body plus all of its call sites simultaneously. For the inner translation unit (right), we translate just the code contained within it. In both cases, we also allow the model to specify any imports it needs. Full prompts are in the Appendix.
3.4 Translation, Validation and Repair
We now attempt to translate each translation unit from unsafe Rust to safe, idiomatic Rust, using an LLM. We pass each unit to an LLM, along with its context slice as described in §˜3.3. We instruct the LLM to convert the translation unit to idiomatic Rust. If the unit is a root translation unit of a function and its context contains call sites, we instruct the LLM to translate the call sites too. We specify a particular format for the LLM response that allows us to easily parse it and separate it into function and call-site translations. If the unit is an inner translation unit, we ask the LLM to ensure that the live-in and live-out variables at the beginning and end of the unit are not affected. This means that translating such a translation unit has no side-effects on any other part of the program. This substantiates our claim in §˜3.2.2 that the order in which we translate the units within a function does not matter.
Sometimes, the translated unit needs additional use statements to work; we allow the LLM to specify this too in its response. An example of a prompt and LLM response is shown in Figure˜4.
We then swap out the original spans of code and swap in the LLM-generated code. We attempt to compile the code and run the test suite. If we encounter an error, we provide it as feedback to the LLM in the same conversation, and ask it to re-generate the entire code, including the function body, call sites and imports. We repeat this for up to $N$ attempts (where $N$ is a parameter of our technique). If one of these $N$ attempts succeeds, we move on to the next function; if all $N$ attempts fail, we restore the original version of the function.
In this manner, we go through the entire program, processing each translation unit exactly once. Throughout this process, we maintain correctness modulo the test cases. Our final program will still have some unsafe functions, but these can be converted to safe Rust by a human expert.
4 Experimental Setup
4.1 Datasets
We contribute a benchmark dataset of 7 C programs collected from GNU coreutils [^2] https://github.com/coreutils/coreutils. These are the basic file, shell and text manipulation utilities of the GNU OS, and include common shell commands like ls and cat. We chose coreutils because it is a large, mature C code project, with many constituent programs of varying sizes. Further, there is an ongoing community effort to manually rewrite coreutils in Rust [^3] https://github.com/uutils/coreutils, which means that a) translating coreutils to Rust is an important real-world problem, and b) future researchers using our benchmark can compare the quality of their automatically-generated idiomatic Rust code with human-written Rust code. Finally, each benchmark in coreutils comes with end-to-end system-level test cases, which have been constructed using fuzzing to ensure high coverage [^4] https://www.pixelbeat.org/docs/coreutils-testing.html.
The 7 programs we picked from coreutils are split, pwd, cat, truncate, uniq, tail, and head. Running C2Rust on these programs was not straightforward, because C2Rust does not work directly with Makefiles of complex programs. Coreutils uses a compilation pattern where some object files are bundled into static libraries, and those are then linked to the necessary executables. This compilation pattern makes it difficult for C2Rust to create a compilation database, which is necessary for further analysis.
To make this problem tractable, we create simpler individual Makefiles for each of the 7 programs, including only the .c and .h files that are strictly necessary for each program. We collect the source files for each program iteratively. Starting with the main C file for that program (e.g.,  split.c for split), we collect the necessary header files and attempt to build the program. If the linking step fails due to missing symbols, we collect the necessary source files that define these symbols and attempt to build the program again. We repeat until the executable is built successfully. The simpler Makefile is then fed to C2Rust.
The statistics of these 7 programs are shown in Table˜I. In addition to these 7 programs from coreutils, we also use benchmarks from a prior work, Laertes [7]. Since evaluating on each additional benchmark incurs a time cost, we selected a subset of 10 out of their 17 benchmarks, which we believe is sufficient for the purposes of a fair evaluation. These benchmarks do not come with end-to-end test cases. Although some of the utilities have unit tests, this is not useful because we cannot run C unit tests on Rust programs.
Program # LoC # Fn
# Cov.
Fn
Max
Fn Len
Mean
Fn Len
Coreutils
split 13848 207 73 1634 52.95
pwd 5859 127 16 902 33.02
cat 7460 166 37 1006 33.84
truncate 7181 124 33 902 44.65
uniq 8299 167 34 902 36.26
tail 14423 266 76 902 43.81
head 8047 153 45 902 40.77
Laertes
tulipindicators 44486 234 306 32.10
xzoom 2524 11 536 134.09
genann 2084 32 201 52.44
optipng 95560 576 42124 143.63
urlparser 1118 22 409 38.41
lil 5400 160 283 30.44
qsort 41 3 17 10.67
snudown 6521 92 792 63.53
grabc 1046 7 61 31.57
bzip2 43374 128 3289 98.84
TABLE I: The statistics of the 7 programs in our coreutils benchmark dataset, plus the 10 programs from Laertes. “# Cov. Fn” is the number of functions covered by tests.
4.2 Baselines
We evaluate against two previous approaches for making C2Rust output more natural.
  • Laertes [7]: Laertes uses a set of simple transformation rules to rewrite each raw pointer as a borrow, and then uses compiler feedback to fix errors until the program compiles.
  • CROWN [30]: CROWN derives ownership constraints for each raw pointer based on their usage patterns, and uses an SMT solver to solve these constraints and transform each raw pointer into a borrow.
4.3 Implementation
We implement C2SaferRust on the nightly-2022-08-08 version of Rust. It is implemented as a cargo subcommand, so it can be run on full Cargo packages. We use the Rust compiler’s High-level Intermediate Representation (HIR) to construct the program call graph and to decompose the program into chunks. The AST traversal in Algorithm˜1 is implemented as an instance of the rustc_hir::intravisit::Visitor trait. The slicing described in §˜3.3 is also performed on the HIR.
We use the Mid-level Intermediate Representation (MIR) to solve the dataflow analysis for each chunk as described in §˜3.3. We use the MIR debug information to find the first and last basic blocks in a chunk, and get the live-in variables of the first block and the live-out variables of the last block. These become the live-in and live-out variables of the chunk.
For all our experiments, the LLM used is gpt4o-mini by OpenAI. We fixed the maximum chunk size, $L$, at 150 lines, and the number of LLM feedback iterations, $N$, at 5. The smallest baseline, qsort, took under 5 mins to translate, and the largest, optipng, took nearly 14 hours. More details about LLM calls and costs are in the Appendix.
4.4 Evaluation
We use two metrics from prior work to evaluate the quality of translation:
  • Raw Pointer Declarations: The number of raw pointer declarations, *const T or *mut T, in the entire program. This includes declarations in function signatures and within structures. This metric was used to evaluate Laertes [7].
  • Raw Pointer Dereferences: The number of raw pointer dereferences throughout the program. This includes all dereferences of the form *<expr> where <expr> evaluates to a raw pointer type (verified by type checking). This metric was also used to evaluate Laertes [7].
Raw pointers are not a feature of idiomatic Rust, and the fewer raw pointers in a program, the more natural it is. Further, dereferencing raw pointers is an unsafe operation which could lead to segmentation faults and other undefined behavior if not used correctly. So for both of the above metrics, a lower value indicates safer code. We further analyze unsafe blocks by investigating their length and types of statements, through the following three metrics:
  • Unsafe Lines of Code: The total lines of code (LOC) contained within all unsafe blocks in the program. This includes the bodies of unsafe functions, which are also treated as unsafe blocks for this metric. A lower value for unsafe LOC indicates a more natural and idiomatic Rust program, as it minimizes the reliance on unsafe constructs.
  • Unsafe Type Casts: Type casts, within unsafe blocks, especially those involving raw pointers (such as ptr as *const u8) are a common pattern which bypass Rust’s strict type system, allowing for potential misuse and may result in undefined behavior if the cast is used incorrectly.
  • Unsafe Call Expressions: Function calls within unsafe blocks, such as std::alloc::alloc(layout) or other custom unsafe functions, allow for a direct interaction with low-level operations. This can be dangerous when encapsulated in unsafe blocks because they bypass Rust’s safety guarantees and potentially lead to memory corruption, dangling pointers, or other undefined behavior.
5 Results
5.1 Effectiveness of C2SaferRust in reducing unsafe constructs (RQ1)
Motivation: In the first research question, we try to evaluate how effectively C2SaferRust can transform C2Rust’s unsafe code into safe and idiomatic Rust.
Evaluation: We run C2SaferRust on each program in our benchmark dataset, which was compiled from coreutils. For each program, we calculated the following metrics both before and after translation (see §˜4.4 for detailed information): (1) the count of raw pointer declarations, (2) the number of raw pointer dereferences, and (3) the extent of unsafe lines of code. Additionally, our analysis was expanded to encompass measurements of (1) unsafe type casts and (2) unsafe call expressions.
TABLE II: Translating CoreUtils to safer Rust using C2SaferRust. Overall results suggest C2SaferRust offers enhanced safety by reducing reliance on raw pointers and unsafe operations.
    Raw Pointer     Raw Pointer     Unsafe     Unsafe Call       Unsafe
    Declarations ($\nabla$)     Deferences ($\nabla$)     Lines ($\nabla$)     Expressions ($\nabla$)     Type Casts ($\nabla$)
Before
After
$\Delta\%$
Before
After
$\Delta\%$
Before
After
$\Delta\%$
Before
After
$\Delta\%$
Before
After
$\Delta\%$
     split     252    214    15    656    540    18    11324    9250    18    2353    2212    6    5979    4663    22
   pwd     164    129    21    295    225    24    4201    3151    25    875    871    0    2248    1563    30
   cat     192    120    38    317    240    24    5625    4189    26    1038    912    12    3116    1992    36
   truncate     156    124    21    326    263    19    5544    4443    20    1040    952    8    3357    2521    25
   uniq     227    166    27    343    250    27    6066    4397    28    1150    1025    11    3590    2340    35
   tail     389    297    24    1092    847    22    11663    8818    24    2580    2433    6    5869    3739    36
   head     192    141    27    442    351    21    6245    4699    25    1378    1189    14    3488    2464    29
     avg. $\Delta\%$             25            22            24            8            30
(a) Our metrics for measuring safety of Rust code, computed across all applications before and after translation.
    Raw Pointer     Raw Pointer     Unsafe     Unsafe Call     Unsafe
    Declarations ($\nabla$)     Deferences ($\nabla$)     Lines ($\nabla$)     Expressions ($\nabla$)     Type Casts ($\nabla$)
Before
After
$\Delta\%$
Before
After
$\Delta\%$
Before
After
$\Delta\%$
Before
After
$\Delta\%$
Before
After
$\Delta\%$
     split     141    130    8    463    405    13    6836    6134    10    1405    1334    5    5400    4427    18
   pwd     38    40    -5    85    74    13    704    565    20    217    223    -3    646    292    55
   cat     93    56    40    242    200    17    3508    $28$    19    659    572    13    2625    1587    40
   truncate     68    66    3    241    211    12    4050    $36$    10    772    676    12    4231    3332    21
   uniq     85    71    16    156    144    8    3027    2636    13    569    516    9    3377    2917    14
   tail     147    125    15    568    475    16    7203    6191    14    1559    1514    3    5619    4129    27
   head     87    75    14    340    297    13    3724    3160    15    903    801    11    3193    2355    26
     avg. $\Delta\%$             13            13            14            7            29
(b) Our metrics for measuring safety of Rust code, computed across all applications, filtered to include covered functions only.
Results: Table˜II tabulates the effectiveness of translating CoreUtils projects into safer Rust using C2SaferRust. In the following, we discuss the key findings in three key metrics: raw pointer declarations, raw pointer dereferences, and line of unsafe code.
  1. 1.
    Raw Pointer Declarations. We observe an average reduction of raw pointer declarations by $\mathbf{25\%}$ with notable reductions across all projects. The most significant improvement is observed in cat, which shows a $\mathbf{38\%}$ reduction in raw pointer declarations. Unexpectedly, we see an increase in this metric for pwd in the filtered seting. This does not signify a decrease in safety of the program; rather, it is due to a quirk of the way this metric is computed. The metric counts only variable bindings as declarations. So foo("abc" as const *u8) would not be counted as a raw pointer declaration, but let arg = "abc" as const *u8; foo(arg) would be.
  2. 2.
    Raw Pointer Dereferences. Unsafe raw pointer dereferences are reduced on average by $\mathbf{22\%}$, demonstrating the tool’s effectiveness in minimizing unsafe manipulations. The project uniq shows the largest decrease at $\mathbf{27\%}$, indicating a significant impact on pointer-heavy codebases.
We also observed a notable reduction in unsafe code across all CoreUtils projects after translation to safer Rust.
  1. 1.
    Lines of unsafe code. Lines of unsafe code decrease by an average of $\mathbf{24\%}$ across all projects, reflecting a significant enhancement in code safety and idiomaticity. uniq shows the largest reduction of $\mathbf{28\%}$, highlighting C2SaferRust’s ability to minimize unsafe code.
  2. 2.
    Unsafe Call Expression: Unsafe function calls decrease, but by a relatively smaller amount compared to the other metrics - an average of $\mathbf{8\%}$ reduction. The most substantial decrease is in head, from 1378 to 1189 casts (by $\approx\mathbf{14}\%$). The sub-par overall performance is due to the reliance on FFI calls, which is discussed further below.
  3. 3.
    Unsafe Type Cast: Unsafe type casts decrease by an average of 30%. The most substantial decrease is in tail, from 5869 to 3739 casts (by $\approx\mathbf{36}\%$).
(a) Original C implementation
(b) C2SaferRust output
(c) C2Rust translation
Figure 5: C code versus C2Rust code versus LLM-translated safer Rust, for the simple_cat function from the cat program.
Discussion: Figure˜5 illustrates a function using the original C code (Fig.˜5), C2Rust translation (Fig.˜5), and C2SaferRust’s safer output (Fig.˜5) for comparison. We make the following observations:
  1. 1)
    C2SaferRust uses idiomatic and native Rust types. Non-native types in the C2Rust code (such as libc::c_char, see 1 in Fig.˜5) are replaced with their native Rust types (i.e., &mut [u8] as seen in 1, Fig.˜5). Note that in some cases these pointers are cast back into their raw pointer-type when passed as an argument to other functions, e.g., 2 and 3 in Fig.˜5.
  2. 2)
    C2SaferRust continues to make use of FFI calls. C2SaferRust accomplishes certain core functionalities through calls to C API functions via foreign function interfaces (FFIs), e.g., safe_read and error (see 2 and 3 in Fig.˜5). This limits the extent of code safety since the arguments and return types of these functions still use raw pointer types. These may only be eliminated when all the external functions are replaced with safe Rust equivalents; this is beyond the scope of this work.
  3. 3)
    C2SaferRust eliminates nonstandard control flow structures and dead code. Rust code generated by the C2Rust transpiler generates a large number of unreachable branches in the C2Rust code. It makes explicit calls to branch conditions that cannot be executed (e.g., if 0 != 0 as seen in 4, Fig.˜5). C2Rust also uses certain macros such as unreachable!() (as seen in 5, Fig.˜5). C2SaferRust effectively eliminates these dead code fragments and produces more concise and idiomatic Rust code.
{myshadowbox}
Summary: C2SaferRust reduces unsafe constructs in Coreutils, cutting raw pointers by 38% and dereferences by 27%. Unsafe lines, type casts and calls decrease by 24%, 8% and 24%, enhancing safety and idiomaticity.
TABLE III: A comparative evaluation of C2SaferRust and other state-of-the-art methods against the baseline (C2Rust). Lower values are preferred. Cells in blue show the greatest improvement on that metric for that project. C2SaferRust presents the most significant improvement as indicated by the best Win/Tie/Loss ratio. Laertes and crown actually cause an increase in the amount of unsafe code in several translations, as indicated by the negative values under the $\%\Delta$ columns.
Raw Pointer
Declaration
Raw Pointer
Dereferences
Lines of
Unsafe Code
Unsafe Call
Expressions
Unsafe
Type Casts
    $\Delta$    $\Delta$    $\Delta$    $\Delta$    $\Delta$
      bzip2     Baseline     227     3764     12594     2580     7614
        Laertes     206 9    3699 2    13115 -4    2788 -8    8699 -14
        crown     194 15    3674 2    12592 0    $29$ -14    7672 -1
        C2SaferRust     199 12    2691 29    10328 18    2455 5    4097 46
      genann     Baseline     73     339     1634     579     1319
        Laertes     73 0    339 0    1650 -1    581 -0    1329 -1
        crown     71 3    317 6    1775 -9    643 -11    1317 0
        C2SaferRust     63 14    273 19    994 39    578 0    362 73
      lil     Baseline     438     1668     4852     1809     2788
        Laertes     430 2    1658 1    4645 4    1844 -2    $27$ 2
        crown     355 19    756 55    5073 -5    3383 -87    2523 10
        C2SaferRust     428 2    1449 13    4582 6    1843 -2    2111 24
      tulipindicators     Baseline     866     1847     17671     2595     12983
        Laertes     866 0    1847 0    25594 -45    2595 0    13050 -1
        crown     865 0    1847 0    17769 -1    2419 7    13195 -2
        C2SaferRust     860 1    1625 12    17314 2    2638 -2    12274 5
      urlparser     Baseline     79     60     845     447     648
        Laertes     74 6    2 97    1002 -19    572 -28    1167 -80
        C2SaferRust     81 -3    51 15    702 17    450 -1    439 32
      grabc     Baseline     13     21     214     48     123
        Laertes     8 38    12 43    213 0    71 -48    122 1
        C2SaferRust     7 46    11 48    84 61    33 31    10 92
      optipng     Baseline     1407     6062     82616     7854     32913
        Laertes     1235 12    5552 8    93696 -13    9189 -17    46450 -41
        C2SaferRust     1197 15    5203 14    64013 23    7354 6    25061 24
      qsort     Baseline     4     10     32     11     12
        Laertes     2 50    6 40    32 0    23 -109    12 0
        C2SaferRust     1 75    6 40    19 41    10 9    13 -8
      snudown     Baseline     244     842     5842     1587     3252
        Laertes     224 8    765 9    7213 -23    1772 -12    5940 -83
        C2SaferRust     231 5    747 11    5583 4    1533 3    2867 12
      xzoom     Baseline     29     172     1468     321     867
        Laertes     29 0    172 0    1530 -4    321 0    897 -3
        C2SaferRust     29 0    168 2    1467 0    325 -1    631 27
Aggregate
Win/Tie/Loss
C2SaferRust     5/1/4     7/1/2     10/0/0     8/0/2     9/0/1
        crown     2/0/8     1/0/9     0/0/10     1/0/9     0/0/10
        Laertes     1/1/8     1/1/8     0/0/10     1/0/9     1/0/9
5.2 Comparison with Existing Tools (RQ2)
Motivation: This section presents a comparative analysis of safe Rust transformations obtained from C2SaferRust, as well as methodologies from contemporary approaches Laertes and crown applied to Laertes benchmark. [^5] The incompatibility of Laertes with the current version of the Rust compiler precluded its execution on our CoreUtils benchmarks due to it being developed on a deprecated compiler version. Therefore, C2SaferRust was instead executed on the 10 Laertes benchmarks666crown and was executed on 4 of the 10 benchmark projects. In the remaining projects, we encountered failure due to limitations in managing unions and variadic arguments [30].
Evaluation: We measure heuristics similar to the previous research question: (1) the number of raw pointer declarations, (2) the number of raw pointer dereferences, and (3) the quantity of unsafe lines of code. Furthermore, we extend our analysis to include measurements of (1) unsafe type casts, (2) unsafe call expressions, and (3) unsafe let declarations.
Results: The results are shown in Table˜III. We observe that C2SaferRust is capable of reducing the size of unsafe blocks, doing better than the baselines in most cases.
  • On the two raw pointer metrics (i.e., raw pointer declarations and raw pointer dereferences), C2SaferRust performs as well as or better than the baselines on a majority of the benchmark programs with an aggregate Win/Tie/Loss of 5/1/4 and 7/1/2 for raw pointer declarations and raw pointer dereferences respectively.
  • C2SaferRust demonstrates superior efficacy in reducing lines of unsafe code, achieving a reduction of up to 61% in grabc. An aggregate Win/Tie/Loss of 10/0/0 shows consistent performance across all applications.
  • We also observed a notable reduction in key unsafe expressions and declarations with C2SaferRust in all projects. On the other hand, Laertes and crown actually cause an increase in unsafe code in several projects, as evidenced by the negative values of $\%\Delta$ in Table˜III under lines of unsafe code, unsafe type casts and unsafe calls.
Section E in the Appendix shows an example of the outputs of C2SaferRust and Laertes, for qualitative comparison.
{myshadowbox}
Summary: C2SaferRust outperforms other approaches with a 39/2/9 aggregate win/tie/loss record over all 5 metrics, enhancing Rust safety and idiomaticity.
5.3 Ablation of C2SaferRust design choices (RQ3)
In this section, we evaluate the individual effect of some of the components of C2SaferRust.
Evaluation: We pick 3 arbitrary baselines from coreutils - cat, uniq and head - and run C2SaferRust in two settings:
  1. 1.
    No Chunking: Translating each function as a whole without dividing it into smaller chunks. Our hypothesis is that LLMs do better when the input is not too long, so we would expect this setting to be worse than the original.
  2. 2.
    Random Order: Translating the functions in a random order, instead of following the ordering described in §˜3.2.2.
Function Length: We also study the effect of function length on translation. For this study, we consider only those functions which are translated as a single unit, i.e., those which are smaller than $L=150$ lines. We define two quantities:
  • Translation Success Rate: The percentage of functions that translate correctly, where a correct translation is one that compiles and passes tests.
  • Number of Attempts: The number of iterations of compiler or test feedback given to the LLM before it generates a successful translation. This is measured only for functions for which a correct translation is generated within $N$ attempts.
We measure Translation Success Rate and Number of Attempts over functions of varying length, from 0 to $L=150$.
(a) Translation Success Rate versus Function Length
(b) #Attempts Required to Translate a Function, vs Length
Figure 6: For our coreutils benchmark dataset - a plot of the fraction of successful translations versus function length, and the number of attempts required to translate a function versus function length. This includes only functions that are translated as a single unit, i.e., functions that are $\leq$ 150 lines.
Results: The results for the component ablation are shown in Table˜IV. There are 3 utilities and 5 metrics, which makes a total of 15 settings. C2SaferRust outperforms both “No Chunking” as well as “Random Order” in 9 out of 15 settings, showing a clear benefit of both design choices.
Figure˜6(a) shows the rate of translation success versus function length, and Figure˜6(b) shows the number of attempts needed to translate a function vs its length. If we ignore the outliers on the right, we can see that as the function length increases, the translation success rate decreases and the number of attempts increases. This provides empirical support for our decision to break up very long functions into smaller chunks for translation.
Program     Method     Raw Ptr     Raw Ptr     Unsafe     Unsafe     Unsafe
        Decls $\downarrow$     Derefs $\downarrow$     Lines $\downarrow$     Calls $\downarrow$     Casts $\downarrow$
cat     Original     192     317     5625     1038     5625
    No Chunking     137     248     3638     922     3638
    Random Order     139     246     4194     955     4194
    Our Tool     120     240     4189     912     4189
uniq     Original     227     343     6066     1150     6066
    No Chunking     169     252     4382     1114     4382
    Random Order     164     258     4516     1117     4516
    Our Tool     166     250     4397     1025     4397
head     Original     192     442     6245     1378     6245
    No Chunking     142     360     4814     1380     4814
    Random Order     142     349     4735     1252     4735
    Our Tool     141     351     4699     1189     4699
TABLE IV: Showing the contribution of the components of our method by ablation on three selected benchmarks.
6 Related Work
Converting C to unsafe Rust. Corrode [5] and C2Rust [9] both aim to convert C to functionally equivalent Rust. C2Rust continues to be actively maintained, whereas Corrode has seen little recent development. Both generate non-idiomatic Rust that relies heavily on unsafe constructs. Citrus [4] similarly provides syntactic mappings from C to Rust, but focuses on being lightweight rather than fully correct: the output may not compile or run without manual revision. It is intended more as a rough starting point than a complete translation pipeline.
Converting unsafe Rust to safer Rust. Because C2Rust code relies extensively on raw pointers, much work has explored automatically replacing them with safer Rust references. Laertes [7] was the first to use compiler feedback to infer lifetimes and mutability, rewriting raw pointers when ownership information could be recovered. Their study also analyzed common patterns of unsafe in C2Rust output. However, the approach applies only when unsafety arises from missing ownership rather than aliasing or other causes, converting about 8.3% of raw pointers. A later extension [8] broadens coverage to more pointer categories and improves conversion rates to about 21%. CROWN [30] also starts from C2Rust output but instead constructs pointer constraints and solves them using an SMT solver. It is more targeted, supporting only mutable, non-array pointers, and converts $\sim$37% of these to safe references. Other efforts pursue narrower correctness improvements: Hong and Ryu [10] replace output-parameter success/failure patterns with Option/Result, and [11] translate C unions into Rust tagged unions. NADER [19] replaces unchecked indexing with checked indexing, allowing a tunable safety–performance tradeoff, but does not address raw pointers or unsafe blocks more generally.
LLMs for Code Translation. Early LLM-based translation work began with TransCoder [21], using Back Translation for fully unsupervised learning. Later, Ahmad et al. [2] enhanced this with code summarization. Roziere et al. [22] generate aligned parallel data for supervised training by generate synthetic pairs using another model and verifying them with unit tests. Recent code-focused LLMs have demonstrated strong translation capabilities [17]. Yang et al. [29] introduce UniTrans, which generates tests and iteratively repairs translations using execution feedback. Nitin et al. [16] use multi-modal specifications with a self-consistency filter, and Yang et al. [28] verify candidate translations against transpiler output using equivalence proofs. However, these works generally deal with programs that fit in a single model context window, such as competitive programming solutions.
Shiraishi and Shinagawa [25] were the first to attempt translation of arbitrarily large repositories, slicing projects into smaller units, translating each with an LLM, then compiling and repairing errors. Their method does not incorporate test feedback and requires each unit to compile independently, limiting scalability. Ibrahimzada et al. [12] translate full Java projects to Python, and Zhang et al. [31] translate Go to Rust with emphasis on avoiding runtime errors. Shetty et al. [24] perform full C→Rust repository translation by dividing code into units, translating with LLMs, and validating using dynamic analysis with LLM-generated tests. Our approach differs in two key ways: (a) we leverage a transpiler to generate Rust skeletons, enabling end-to-end correctness checking at every iteration; and (b) we decompose code below method granularity, allowing translation of repositories with extremely large functions.
7 Threats to Validity
Although C2Rust makes efforts to ensure functional equivalence between the C and unsafe Rust code, this is not a formal guarantee. C2SaferRust uses C2Rust as the first step, and so the validity of the entire approach depends on the correctness of C2Rust’s transpilation. However, C2Rust merely provides an interface between the original code and translated code that enables the program to be tested end-to-end. It could be replaced with any other such framework, like Rust’s Foreign Function Interface (FFI).
Our metrics are intended to model the safety and naturalness of the code; however it is possible that an improvement in a metric doesn’t correlate with improved safety or naturalness of the code. For example, the number of lines of code within unsafe blocks could decrease simply by eliminating line breaks and using semicolons to delimit lines. We therefore assume that the tools we consider do not attempt to game the metrics by performing such manipulations.
The code we generate still uses a significant amount of unsafe code. Some of this is because we do not translate code outside functions, like structure definitions, and interacting with these structures could necessitate the use of some unsafe code. Another observation is that the LLM does not replace C FFI function calls with their safe Rust equivalents; rather, it uses unsafe code to wrap these FFI calls.
8 Conclusion
In this work, we have presented C2SaferRust, a system to automatically translate full C repositories to safer Rust. Based on the novel insight that the C2Rust transpiler can be used as an intermediary to bridge between C and Rust, we first convert the entire repository into unsafe Rust. We divide the code repository into smaller “translation units”, and extract each translation unit along with its context information. We use an LLM to translate each unit to safer Rust, swap in this translation, and propagate the changes to the rest of the code base. We contribute a dataset of 7 C programs from GNU CoreUtils with their corresponding C2Rust transpiled versions. Each program comes with end-to-end test cases, allowing one to empirically verify the correctness of translation. On this dataset, we show that we reduce raw pointer declarations and dereferences by up to 38%, and reduce unsafe code by up to 28%. We also compare our tool with two baselines for making Rust code safer, and observe consistent gains. Future work could look at translating code outside functions, like structure definitions, as well as remapping C FFI calls to safe Rust equivalents.
Appendix A Decomposition Algorithm
1input: G - the AST of the program
2input: L - the max length of a translation unit
3
4output: units - a list of pieces of code, each
5 having $\leq$ L lines
6
7units := []
8walk_subtree(G)
9
10fn walk_subtree(node):
11 for child in node.children:
12 match child.type:
13 Func => visit_func(child),
14 Block => visit_block(child),
15 _ => walk_subtree(child)
16
17fn visit_func(func):
18 // Traverse each function longer than L lines
19 if line_count(func) <= L:
20 chunks.push(func)
21 else:
22 walk_subtree(func)
23
24// A block is any entity within {..} parantheses
25fn visit_block(block):
26 // Walk each subtree longer than L lines
27 for stmt in block.statements:
28 if line_count(stmt) > L:
29 walk_subtree(stmt)
30 lower, upper := 0, 0 // Maintain two cursors
31 while upper << block.statements.len():
32 if line_count(block.statements[lower..upper]) > L:
33 // Adding this statement caused the
34 // length to spill over
35 unit := block.statements[lower..upper-1]
36 units.push(unit)
37 // Remove this unit from the AST
38 G.prune_nodes(unit)
39 lower = upper
40 else:
41 upper += 1
42 // We have some leftover lines
43 if lower << block.statements.len():
44 unit := block.statements[lower..]
45 units.push(unit)
46 // Remove this unit from the AST
47 G.prune_nodes(unit)
Algorithm 1 Decomposing the program into translation units, each having fewer than $L$ lines.
Appendix B More Coreutils Results
In this section, we present some more results on the Coreutils datasets, adding to the results in Table˜II. Table˜V contains the results of a variant of our approach that uses direct C to Rust translation. Table˜VI shows the number of LLM calls and approximate usage cost for all our benchmarks.
TABLE V: Translating CoreUtils to safer Rust using a variant of C2SaferRust, with direct C to Rust translation.
    Raw Pointer     Raw Pointer
    Declarations ($\nabla$)     Deferences ($\nabla$)
Before
After
$\Delta\%$
Before
After
$\Delta\%$
     split     252    232    8    656    628    4
   pwd     164    138    16    295    265    10
   cat     192    163    15    317    289    9
   truncate     156    142    9    326    278    15
   uniq     227    179    21    343    310    10
   tail     389    338    13    1092    1027    6
   head     192    172    10    442    420    5
(a) Number of raw pointer declarations and dereferences.
    Unsafe     Unsafe     Unsafe Call
    Lines ($\nabla$)     Type Casts ($\nabla$)     Expressions ($\nabla$)
Before
After
$\Delta\%$
Before
After
$\Delta\%$
Before
After
$\Delta\%$
     split     11324    10357    9    2353    2326    1    11324    10357    9
   pwd     4201    2921    30    875    851    3    4201    2921    30
   cat     5625    4663    17    1038    950    8    5625    4663    17
   truncate     5544    4581    17    1040    1050    -1    5544    4581    17
   uniq     6066    4986    18    1150    1129    2    6066    4986    18
   tail     11663    9780    16    2580    2485    4    11663    9780    16
   head     6245    5552    11    1378    1383    -0    6245    5552    11
(b) Number of unsafe expressions and declarations.
Program # Fn
# Trans.
Units
LLM
Calls
Approx
Cost (\$)
Coreutils
split 207 292 1140 0.68
pwd 127 152 527 0.31
cat 166 205 636 0.38
truncate 124 170 599 0.36
uniq 167 213 776 0.46
tail 266 350 1224 0.73
head 153 193 666 0.40
Laertes
tulipindicators 234 242 1017 0.61
xzoom 11 26 101 0.06
genann 32 35 124 0.07
optipng 576 941 3518 2.10
urlparser 22 25 101 0.06
lil 160 184 800 0.48
qsort 3 3 12 0.01
snudown 92 108 488 0.29
grabc 7 7 22 0.01
bzip2 128 251 885 0.53
TABLE VI: LLM usage cost estimates for translating each benchmark using C2SaferRust with gpt4o-mini.
Appendix C Prompts
Here is a function:
‘‘‘rust
unsafe extern "C" fn gettext_quote(
mut msgid: *const libc::c_char,
mut s: quoting_style,
) -> *const libc::c_char {
let mut translation: *const libc::c_char = gettext(msgid);
let mut locale_code: *const libc::c_char = 0 as *const libc::c_char;
if translation != msgid {
return translation;
locale_code = locale_charset();
Here are its call sites
Call site 1:
‘‘‘rust
right_quote = gettext_quote(
b"’\0" as *const u8 as *const libc::c_char,
quoting_style,
Call site 2:
‘‘‘rust
left_quote = gettext_quote(
b"‘\0" as *const u8 as *const libc::c_char,
quoting_style,
The file contains the following imports:
‘‘‘rust
use ::libc;
Convert the function to idiomatic Rust, meaning Rust code that does not make use of features like unsafe, raw pointers, and the C API whenever possible. Do not change the function name.
Follow the following format for your output: Place the function translation between the tags <FUNC> and </FUNC>. If the function signature changed in translation, its callsites will need to be modified as well. Place each callsite translation (in the same order it appears above) between <CALL> and </CALL>. Note that even if the callsite is only a single statement, the translation can be mutiple statements. For example, you may need to declare new variables, or convert between types, either before or after the call. The translation should be such that the surrounding code is not affected by the changes.
Any functions or variables without definitions are defined elsewhere in the code. Do not attempt to redefine them or import them.
If you are using any new functions and you need new imports for those, place them between the tags <IMPORTS> and </IMPORTS>. This should be only *new* imports. Do not include existing imports.
DO NOT include markdown characters like "‘‘‘" or "‘‘‘rust" in your translation.
TABLE VII: Translation prompt for a Coreutils function.
Appendix D Test Statistics
Table˜VIII shows the statistics of the Coreutils test suite that we used to evaluate the correctness of our translations. Each test is a Bash or Perl script that is run against the compiled executable.
Program # Fn
# Cov.
Fn
# Test
Scripts
Coreutils
split 207 73 12
pwd 127 16 2
cat 166 37 4
truncate 124 33 8
uniq 167 34 3
tail 266 76 30
head 153 45 4
TABLE VIII: The statistics of the test suite for the 7 programs in our coreutils benchmark dataset. “# Cov. Fn” is the number of functions covered by tests, and “# Test Scripts” is the number of end-to-end test scripts.
Appendix E More Qualitative Comparisons
In this section, we show some more examples of C2SaferRust output that will enable us to qualitatively analyze its performance.
Figure˜7 shows an example of the output C2Rust vs Laertes vs C2SaferRust. This is taken from qsort, one of the Laertes benchmarks. We can see that although both Laertes and C2SaferRust eliminate the raw pointer *mut, the output of C2SaferRust is much more natural and easy to follow.
pub unsafe extern "C" fn swap(
mut a: *mut std::os::raw::c_int,
mut b: *mut std::os::raw::c_int) {
let mut t: std::os::raw::c_int = *a;
*a = *b;
*b = t;
(a) C2Rust output
pub unsafe extern "C" fn swap<’a1, a2>(
mut a: Option<&’a1 mut std::os::raw::c_int>,
mut b: Option<&’a2 mut std::os::raw::c_int>) {
let mut t: i32 = *(borrow_mut(&mut a)).unwrap();
*(borrow_mut(&mut a)).unwrap() =
*(borrow_mut(&mut b)).unwrap();
*(borrow_mut(&mut b)).unwrap() = t;
(b) Laertes output.
pub fn swap(a: &mut i32, b: &mut i32) {
let t = *a;
*a = *b;
*b = t;
(c) C2SaferRust output.
Figure 7: A qualitative comparison of Laertes and C2SaferRust.
Figure˜8 shows the function ireallocarray, which is a common helper function used by all the Coreutils programs. It takes a pointer pointing to potentially uninitialized memory, and allocates memory for it such that the total allocation is equal to a certain specified number of bytes. This is implemented in C with a void * pointer. The C2Rust code uses *mut libc::c_void, which is not an idiomatic Rust type. On the other hand, C2SaferRust correctly handles this by using the type Option<Vec<u8>> for p. This also allows for the possibility that the memory might be uninitalized, which would correspond to the None variant of the Option enum.
/* ireallocarray (ptr, num, size) is like reallocarray (ptr, num, size).
It returns a non-NULL pointer to num * size bytes of memory.
Upon failure, it returns NULL with errno set. */
IALLOC_INLINE void *
ireallocarray (void *p, idx_t n, idx_t s)
if (n <= SIZE_MAX && s <= SIZE_MAX)
size_t nx = n;
size_t sx = s;
if (n == 0 $|$ s == 0)
nx = sx = 1;
p = reallocarray (p, nx, sx);
return p;
else
return _gl_alloc_nomem ();
(a) C code
pub unsafe extern "C" fn ireallocarray(
mut p: *mut libc::c_void,
mut n: idx_t,
mut s: idx_t,
) -> *mut libc::c_void {
if n as libc::c_ulong <= 18446744073709551615
as libc::c_ulong
&& s as libc::c_ulong <= 18446744073709551615
as libc::c_ulong
let mut nx: size_t = n as size_t;
let mut sx: size_t = s as size_t;
if n == 0 as libc::c_int as libc::c_long
$|$ s == 0 as libc::c_int as libc::c_long
sx = 1 as libc::c_int as size_t;
nx = sx;
p = reallocarray(p, nx, sx);
return p;
} else {
return _gl_alloc_nomem()
(b) C2Rust output.
pub fn ireallocarray(
p: Option<Vec<u8>>,
n: idx_t,
s: idx_t,
) -> Option<Vec<u8>> {
if n as usize <= usize::MAX && s as usize <= usize::MAX {
let mut nx: usize = n as usize;
let mut sx: usize = s as usize;
if n == 0 $|$ s == 0 {
sx = 1;
nx = sx;
let mut vec = p.unwrap_or_else($|$
Vec::with_capacity(nx * sx)
vec.resize(nx * sx, 0);
return Some(vec);
} else {
return None;
(c) C2SaferRust output.
Figure 8: A comparison between C, C2Rust, and C2SaferRust, for an example involving memory allocation. C2SaferRust appropriately handles the allocation using Option<Vec<u8>> in Rust.